This question is from quadratic equations. In which of these cases will the roots of ax^2+bx+c=0 ,a is not equal to zero; a,b,c belong to real numbers will be definitely real. a) a (a+b+c)0 c) c (4a-2b+c)

Grade:11

1 Answers

Riddhish Bhalodia
7 years ago
In the case of
c(4a+2b+c) = f(0)f(2) = 0
so either one of the two will be zero and hence a real root, thus in the case when c(4a+2b+c) = 0 we will definitely have real roots